A AND B ARE TWO IDENTICAL VESSELS.A CONTAINS 15g ethane and b contains gas x2 at 298k and 1 atm respectively.vapour density of gas x2?

Grade:Select Grade

2 Answers

Rohit Mamoria
7 years ago
see pv =nrt here p, v and t are constant therefore n is constant now the ratio of given mass and molecular mass also will be constant 
therfore
15/30 = 75/molecular mass of gas 
which gives the unknown molecular mass as 150
now we know vapur density is molecular mass /2
therefore 150/2 or 75
and 75 is the answer 
Sagar
4 years ago
Given : for Ethane  ;
Wt.of ethane = W1 = 15g , Mol. Wt. = M1 = 30
for 'X' ;
Wt. Of 'X' = 75g , Mol. Wt. = M2 = ?
By using Avogadro's Law - (volume -mole relationship) :
»»  V1/n1 = V2/n2
But,  for constant volume, 
»»  n1 = n2
»»  W1/M1 = W2/M2
»»  15/30 = 75/M2
.°.   M2 = 150 g/mol
Now, we have, 
Vapour density = Molecular Wt. /2
.°. Vapour density of 'X' = M2/2 = 150/2 = 75
Hence,  Vapour density of 'X' =  75  ,,
 
✨Hope it helps ✨